4x + 2(40-x) = 100
1
Expert's answer
2013-05-14T09:18:20-0400
4x + 2(40-x) = 100
At first, let's open the braces:
4x + 80-2x = 100
Now let's gather terms containing x in the left part, others in the right:
4x -2x = 100-80
Simplifying the equation we get:
2x = 20& whence x=10

Let's check the obtained answer:
4*10 + 2(40-10) = 100
40+60=100 that's right

Answer: x=10